1. According to the following reaction, how many grams of iron(III) oxide are required for the complete reaction of 28.0 grams of hydrochloric acid?

hydrochloric acid (aq) + iron(III) oxide (s) water (l) + iron(III) chloride (aq)


2. According to the following reaction, how many grams of hydrobromic acid are needed to form 23.4 grams of bromine?

hydrobromic acid (aq) hydrogen (g) + bromine (l)
1
Expert's answer
2019-10-03T05:19:25-0400

1) First we should equalize the equation:


6HCl + Fe2O3 = 2FeCl3 + 3H2O


Now, we should found the amount in moles of HCl and that of Fe2O3 would be 6 times less:


"n(HCl)=\frac{28.0g}{36.5g/mol}=0.767mol"


"n(Fe_2O_3)=\frac{1}{6}*0.767mol=0.128mol"


Thus, the mass of Fe2O3 is equal to:


"m(Fe_2O_3)=160g\/mol*0.128mol=20.5g"


2) Lets equalize it:


2HBr = H2 + Br2


Now, lets calculate the amount of bromine in moles and the amount of HBr would be 2 times Br2 amount:


"n(Br_2)=\frac{23.4g}{160g/mol}=0.146mol"


"n(HBr)=2*0.146mol=0.292mol"


Now, it is the right time to calculate the mass of HBr:


"m(HBr)=81g\/mol*0.292mol=23.7g"


Thus, the mass of HBr required is equal to 23.7g.